OpenStudy (anonymous):

A private school opened in 2005 with an initial enrollment of 85 students. The enrollment has increased by an average of 18 students each year since the school opened. A nearby public school had an enrollment of 95 in 2005, and its enrollment has increased by an average of 15 students per year. Let t = time in years, with t = 0 representing the year 2005. Let f(t) = the number of students enrolled at the private school and g(t) = the number of students enrolled at the public school. Create the two functions to represent the situation. Graph the system of equations to determine the the point of intersection. Interpret the point of intersection in the context of the problem.

OpenStudy (anonymous):

http://flockdraw.com/8cq4mv

OpenStudy (anonymous):

What is with the link?

OpenStudy (anonymous):

its so i we can work it out together

OpenStudy (anonymous):

Do you know how to draw a line?

OpenStudy (anonymous):

? yes i dont how to do this prblem

OpenStudy (anonymous):

First think is you want to list out the value of each function for different values of \(t\).

OpenStudy (anonymous):

? ><ok

OpenStudy (anonymous):

"2005 with an initial enrollment of 85 students. The enrollment has increased by an average of 18 students each year since the school opened"

OpenStudy (anonymous):

"2005 with an initial enrollment of 85 students" We'll say that \(t=0\) at 2005 since that is where we start to keep track.

OpenStudy (anonymous):

Our function will look like: \[ f(t) = mt+b \]where \(m\) is slope and \(b\) is \(y\) intercept. Do you follow?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

This is the equation of a line.

OpenStudy (anonymous):

The slope intercept form.

OpenStudy (anonymous):

mhh o.o

OpenStudy (anonymous):

They never taught you about slope intercept form?

OpenStudy (anonymous):

y=mx+b?

OpenStudy (anonymous):

Yes, but this time the input variable is \(t\) not \(x\). And the output is \(f(t)\) not \(y\)

OpenStudy (anonymous):

Other than that, everything is the same.

OpenStudy (anonymous):

initial enrollment of 85 This means \(b=85\) increased by an average of 18 This means \(m=18\)

OpenStudy (anonymous):

mhh o.o i c

OpenStudy (anonymous):

So the first equation is \[ f(t)=18t+85 \]

OpenStudy (anonymous):

Can you do the same thing for: " A nearby public school had an enrollment of 95 in 2005, and its enrollment has increased by an average of 15 students per year."

OpenStudy (anonymous):

yes o.o

OpenStudy (anonymous):

Okay let's see what you get.

OpenStudy (anonymous):

f(t)=mt+b f(t)=15t +95?

OpenStudy (anonymous):

". Let f(t) = the number of students enrolled at the private school and g(t) = the number of students enrolled at the public school." So this time we want to use \(g(t)\) not \(f(t)\)

OpenStudy (anonymous):

so its still the same ? but juss with a g?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

g(t)=15t +95 i c ! : )

OpenStudy (anonymous):

Now you want to see what happen when \(t=0,1,2,3\)

OpenStudy (anonymous):

That will give you some points.

OpenStudy (anonymous):

can we do it on the website ? i also need to show the graph >~<

OpenStudy (anonymous):

ok plot the points.

OpenStudy (anonymous):

so i substitute for t =0,1,2,3,

OpenStudy (anonymous):

does the website still work ?

OpenStudy (anonymous):

Yes, then you can plot the points.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

okay : ) so f(0)=18(0)+85

OpenStudy (anonymous):

how would or wat do i need to do now?

OpenStudy (anonymous):

f=0+85

OpenStudy (anonymous):

\[f(0) = 18(0)+85=0+85=85\]

OpenStudy (anonymous):

So one point is \((0,85)\)

OpenStudy (anonymous):

http://flockdraw.com/8ztk5e

OpenStudy (anonymous):

do i need a legend or key?

OpenStudy (anonymous):

ON the bottom you should put like 0,1,2,3... On the side you should put maybe 80,90,100,110

OpenStudy (anonymous):

ookay i did that

OpenStudy (anonymous):

so f(1)=18(1)+85 --> f=18+85 f= 100 but i plot it as 18, 85?

OpenStudy (anonymous):

The point would be \((1,100)\)

OpenStudy (anonymous):

Each point is \((t,f(t))\)

OpenStudy (anonymous):

O-O ohh so.. if i do f(3)=18(3)+85 --> F=54+85 F=139 . so it would be (1,139)?

OpenStudy (anonymous):

ohh (3,139)?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Now you can find points for \(g(t)\)

OpenStudy (anonymous):

yes ! : ) so i do the same right ? to the g(t)

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

students is y and /top x is years ? bottom

OpenStudy (anonymous):

so for F ive got 0,85 1.103 2,121 3,139 G 0,95 1,110 2,125 3,140

OpenStudy (anonymous):

Okay, good. Now you need to plot those points on a graph and you're done.

OpenStudy (anonymous):

do i plot G like i do to F?

OpenStudy (anonymous):

Yes, plot them on the same graph.

OpenStudy (anonymous):

You might also want to plot \(t=4\) as well.

OpenStudy (anonymous):

wat u mean plot t=4?

OpenStudy (anonymous):

Meaning plot \((4,f(4))\) and \((4,g(4))\)

OpenStudy (anonymous):

4,4 for both G and F?

OpenStudy (anonymous):

No, what is \(f(4)\)?

OpenStudy (anonymous):

f(4)=18(4)+85?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

oh ! : ) okay = (4,157

OpenStudy (anonymous):

Now do that for g(t) and we'll have enough points.

OpenStudy (anonymous):

4,155

OpenStudy (anonymous):

Yeah, just plot all the points and you'll get it.

OpenStudy (anonymous):

http://flockdraw.com/8ztk5e

OpenStudy (anonymous):

Now use line tool

OpenStudy (anonymous):

?

OpenStudy (anonymous):

That shows you what the lines would look like.

OpenStudy (anonymous):

i tho i plot 0,85 1.103 2,121 3,139 & 0,95 1,110 2,125 3,140 ?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

interception is ? 4 ?

OpenStudy (anonymous):

@wio ? ><

OpenStudy (anonymous):

It is near 4